OpenStudy (anonymous):

it takes joaquin 1/5 of an hour to hike 1/4 mile what is joaquins unit rate in miles per hour HELP!!!

OpenStudy (texaschic101):

(1/5) / (1/4) = x / (4/4) -- 1/5 hr to 1/4 mile = x hr to 4/4 mile or 1 mile) cross multiply (1/4)(x) = (1/5)(4/4) 1/4x = 1/5 x = (1/5) / (1/4) x = 1/5 * 4/1 x = 4/5 hr so the unit rate per mile is (4/5) / 1 or 4/5 of a mile per hr

OpenStudy (anonymous):

im confused!!